# Regime Learning for Differentiable Particle Filters

### v1

@@ III-A Regime switching particle filtering
6:     Sample $\tilde{\mathbf{x}}^{i}_{t}=Q\left({\tilde{\mathbf{x}}^{i}_{t}|\tilde{\mathbf{x}}^{A^{i}_{t}}_{t-1}}\right)$;
@@
$\hat{\mathbf{x}}_{t}=\sum^{N}_{i=0}\tilde{\mathbf{x}}^{i}_{t}\bar{w}^{i}_{t},$ (2)
@@
$\hat{p}\left({\mathbf{y}_{0:t}}\right)=\prod^{t}_{s=0}\frac{1}{N}\sum^{N}_{i=0}w^{i}_{t}.$ (3)
@@ IV-C Training the RLPF
For the supervised term, we run a particle filter on an SSM defined as in Section IV-A, with the state as $\left\{{\mathbf{x}_{t},k_{t},\mathbf{r}_{t}}\right\}$ and the observation as $\mathbf{y}_{t}$. For the unsupervised term, we treat all the information we have access to in training as observed, i.e. take $\left\{{\mathbf{x}_{t},\mathbf{y}_{t}}\right\}$ to be the observations. However, in doing so the observations depend on their history not only through the state, so is not, by our definition, an SSM. However, treating the past observations as constant, it is easy to show that the usual particle filtering algorithm remains correct. We refer to this strategy as a “marginal filter” as it runs a particle filter only on the marginal process that is the switching dynamic.
@@
The intuition behind this strategy is that training on the ELBO loss minimises a KL-divergence to bring the learned model closer to the true model; and the MSE loss serves to guide the training towards a lower validation objective. We present the full RLPF procedure in Algorithm 2.111Our implementation, that was used to generate all results reported in this paper, can be found at: https://github.com/John-JoB/Regime_Switching
@@
5:     Run a particle filter (Algorithm 1) to obtain $w^{1:N}_{0:T}$ with inputs: $M_{0}\leftarrow Q^{\theta}_{0}\bullet R^{\theta}$,       $Q_{0}\leftarrow\text{Uniform}\left({k\in\mathcal{K}}\right)\bullet R^{\theta}$ $M\leftarrow K^{\theta}\bullet R^{\theta}$        $Q\leftarrow\text{Uniform}\left({k\in\mathcal{K}}\right)\bullet R^{\theta}$ $G_{0}\leftarrow G^{\theta}\bullet M^{\theta}_{0}$       $G_{t\geq 1}\leftarrow G^{\theta}\bullet M^{\theta}\left({\cdot|\mathbf{x}_{t-1}}\right)$$T\leftarrow T$                  $N\leftarrow N$ $\mathbf{y}_{0:T}\leftarrow\left\{{\mathbf{y}_{0:T},\mathbf{x}_{0:T}}\right\}$
@@ V-C Baseline Models
Both the RSDBPF and RSPF require access to stronger priors on the model than the proposed setting allows and are included as oracle algorithms.
@@ V-E Results
We present the results for both experiments in Table I including baselines and ablation results. For the RLPF and the RSDBPF the qualification -$\lambda$ or -MSE denotes whether the experiment used the proposed training algorithm (Section IV-C), or the classical MSE only approach.
For both experiments, the proposed RLPF-$\lambda$ leads to the smallest mean squared errors among non-oracle algorithms. Moreover, being a particle filter, it has a statistical interpretation; with it one can evaluate properties such as uncertainties, the data likelihood, and make predictions about future state or data. The LSTM is the next best in the reported error metrics, but it does not offer the statistical properties seen in particle filters.
We also note that for both the RSDBPF and the RLPF, the variants trained using the proposed training objective achieved lower MSEs and variances than the respective pure-MSE approaches.

### v2

@@ III-A Regime switching particle filtering
6:     Sample $\tilde{\mathbf{x}}^{i}_{t}\sim Q\left({\tilde{\mathbf{x}}^{i}_{t}|\tilde{\mathbf{x}}^{A^{i}_{t}}_{t-1}}\right)$;
@@
$\hat{\mathbf{x}}_{t}=\sum^{N}_{i=1}\tilde{\mathbf{x}}^{i}_{t}\bar{w}^{i}_{t},$ (2)
@@
$\hat{p}\left({\mathbf{y}_{0:t}}\right)=\prod^{t}_{s=0}\frac{1}{N}\sum^{N}_{i=1}w^{i}_{t}.$ (3)
@@ IV-C Training the RLPF
For the supervised term, we run a particle filter on an SSM defined as in Section IV-A, with the state as $\left\{{\mathbf{x}_{t},k_{t},\mathbf{r}_{t}}\right\}$ and the observation as $\mathbf{y}_{t}$. For the unsupervised term, we treat all the information we have access to in training as observed, i.e. take $\left\{{\mathbf{x}_{t},\mathbf{y}_{t}}\right\}$ to be the observations. However, in doing so the observations depend on their history not only through the state, so is not, by our definition, an SSM. However, treating the past observations as constant, it is easy to show that the usual particle filtering algorithm remains correct. We refer to this strategy as a “marginal filter” as it runs a particle filter only on the marginal process that is the switching dynamic. The intuition behind this strategy is that training on the ELBO loss maximises the data likelihood to bring the learned model closer to the true model; and the MSE loss serves to guide the training towards a lower validation objective.
@@
We present the full RLPF procedure in Algorithm 2.111Our implementation, that was used to generate all results reported in this paper, can be found at: https://github.com/John-JoB/Regime_Switching
@@
5:     Run a particle filter (Algorithm 1) to obtain $w^{1:N}_{0:T}$ with inputs: $M_{0}\leftarrow K^{\theta}_{0}\bullet R^{\theta}$,       $Q_{0}\leftarrow\text{Uniform}\left({k\in\mathcal{K}}\right)\bullet R^{\theta}$ $M\leftarrow K^{\theta}\bullet R^{\theta}$        $Q\leftarrow\text{Uniform}\left({k\in\mathcal{K}}\right)\bullet R^{\theta}$ $G_{0}\leftarrow G^{\theta}\bullet M^{\theta}_{0}$       $G_{t\geq 1}\leftarrow G^{\theta}\bullet M^{\theta}\left({\cdot|\mathbf{x}_{t-1}}\right)$$T\leftarrow T$                  $N\leftarrow N$ $\mathbf{y}_{0:T}\leftarrow\left\{{\mathbf{y}_{0:T},\mathbf{x}_{0:T}}\right\}$
@@ V-C Baseline Models
We include two oracle approaches: the RSDBPF which simulates from the true switching dynamic but learns the individual regimes; and the RSPF which corresponds to a particle filter run on the true model.
@@ V-E Results
We present the results for both experiments in Table I including baselines and ablation results, and a comparison of the computation costs under the Pólya-Urn experiment in Table II for the non-oracle methods. The Markov times were similar so are excluded for brevity. For the RLPF and the RSDBPF the qualification -$\lambda$ or -MSE denotes whether the experiment used the proposed training algorithm (Section IV-C), or the classical MSE only approach. All timing experiments were performed using an NVIDA RTX 4090 GPU.
For both experiments, the proposed RLPF-$\lambda$ leads to the smallest mean squared errors among non-oracle algorithms. Moreover, being a particle filter, it has a statistical interpretation; with it one can evaluate properties such as uncertainties, the data likelihood, and make predictions about future state or data. The LSTM is the next best in the reported error metrics, but it does not offer the statistical properties seen in particle filters. We also note that for both the RSDBPF and the RLPF, the variants trained using the proposed training objective achieved lower MSEs and variances than the respective pure-MSE approaches.
Aside from the Transformer, all algorithms considered scale linearly in time with trajectory length. However, the relatively simple architecture of the LSTM and better ability to take advantage of GPU parallelism give it a significant speed advantage, as shown in Table II. Achieving better parallelism for the RLPF to bring its time cost closer to that of the DPF is left for future work.
@@
TABLE II: Average computation times per training epoch (10 batches of 100 parallel filters of 200 particles each), and testing run (1 batch of 500 parallel filters of 2000 particles each) on the Pólya experiment.
Algorithm Av. train epoch time (s) Av. test time (s)
Transformer (baseline) $0.182$ $0.00310$
LSTM (baseline) $\mathbf{0.0145}$ $\mathbf{0.000792}$
DBPF (baseline) $0.553$ $0.243$
MADPF (baseline) $18.6$ $3.73$
RLPF-MSE (baseline) $4.42$ $0.620$
RLPF-$\lambda$ (proposed) $6.66$ $0.612$